# Fine-Tuning a Simulation-Driven Estimator

### v2

@@ Abstract
I Introduction
Recent advances in high-fidelity simulation have popularized digital twins (DTs)—software replicas of physical systems that let engineers explore operating conditions safely and cheaply before deployment [8, 9, 15]. In practice, once a DT is engineered, most modeling choices are fixed; what remains is calibrating a small set of parameters so the DT replicates a specific physical system.
When a DT is available, one can generate synthetic input–output data [14] over a range of parameter settings and train a supervised machine learning model that maps observations to these settings, thereby yielding a simulation-driven estimator. At implementation, this simulation-driven estimator takes real measurements and returns parameter estimates [2, 5, 6]. The Two-Stage (TS) estimator [4, 5] is an attractive simulation-driven estimator with a simple structure that involves two stages: (i) first, the observations are compressed into representative features, and then (ii) these compressed features, along with their corresponding parameter values, are used to train a supervised machine learning model. This structure avoids the need for an explicit likelihood function and it requires low computation time for producing estimates, due to extensive offline training, compared to standard point estimators like the prediction error method (PEM) and the dual Extended Kalman Filter (dual-EKF) [10, App. B.2]. The TS estimator can also be used to learn control design rules, effectively taking the human out of the loop [1].
@@ I Introduction
Our novel fine-tuning procedure first compares the compressed features from real data to those produced by the DT at the TS initial prediction, to determine whether an OOD case has occurred by applying a statistical test based on a whitened feature discrepancy measure constructed from estimated feature covariances. If an OOD case is detected, we: (i) compute Gauss–Newton (GN) iterations on the initial prediction by TS, (ii) generate a confidence set around the GN update and re-simulate a targeted local training dataset with the DT, and (iii) fine-tune the final portion of the second stage of the TS estimator, which is inspired from transfer learning [13] approaches to system identification [3, 12]. In particular, our contributions are the following:
@@
The remainder of the paper is organized as follows. In Section II, we introduce the problem statement. In Section III, our method for fine-tuning the pretrained TS estimator is detailed, and in Section IV, we demonstrate the effectiveness of our approach through several numerical examples. Finally, Section V concludes the paper.
II Problem Statement
We consider a data-generating mechanism $\mathcal{M}(\boldsymbol{\theta})$—a high-fidelity simulator or digital twin (DT)—parameterized by $\boldsymbol{\theta}\in\Theta\subseteq\mathbb{R}^{d}$. For notational simplicity, we assume a single-input single-output (SISO) setting: given an $N$-length scalar input sequence $\boldsymbol{u}=(u_{1},\dots,u_{N})^{\top}\in\mathbb{R}^{N}$, the DT produces a (possibly stochastic) scalar output sequence $\boldsymbol{y}=(y_{1},\ldots,y_{N})^{\top}\in\mathbb{R}^{N}$. Concretely, $\boldsymbol{y}=\boldsymbol{y}(\boldsymbol{\theta};\omega)$, where $\omega$ denotes the simulator’s noise seed (or any internal randomness). Assume that a length-$N$ trajectory $\boldsymbol{z}_{0}=(u_{1},y_{0,1},\ldots,u_{N},y_{0,N})\in\mathbb{R}^{2N}$ has been measured from the DT at some unknown parameter $\boldsymbol{\theta}_{0}\in\Theta$. The goal is to estimate $\boldsymbol{\theta}_{0}$ from $\boldsymbol{z}_{0}$ by constructing a simulation-driven parameter estimator.
Given a calibration set $\Theta_{p}\subseteq\Theta$, constructed from prior domain knowledge and anticipated operating regimes, we draw i.i.d. parameter settings $\{\tilde{\boldsymbol{\theta}}_{i}\}_{i=1}^{m}$ from $\Theta_{p}$. For each draw $\tilde{\boldsymbol{\theta}}_{i}$, we run the simulator with (possibly) its own noise seed $\omega_{i}$ and the chosen input sequence [^1] Due to space limitations, in this paper the input sequence is considered fixed, even though the treatment can be extended to random inputs. $\boldsymbol{u}$, obtaining $\boldsymbol{z}_{i}=(u_{1},y_{i,1},\ldots,u_{N},y_{i,N})$. This yields a synthetic, labeled dataset $\mathcal{D}=\{(\boldsymbol{z}_{i},\tilde{\boldsymbol{\theta}}_{i})\}_{i=1}^{m}$. An estimator $\hat{\theta}_{\mathrm{pre}}\colon\mathbb{R}^{2N}\!\to\mathbb{R}^{d}$ is then fitted by solving the following minimization problem over a prescribed function class $\mathcal{F}$ (e.g. a neural network):
$\hat{\theta}_{\mathrm{pre}}(\cdot)\in\arg\min_{F\in\mathcal{F}}\frac{1}{m}\sum_{i=1}^{m}L\!\left(F(\boldsymbol{z}_{i}),\,\tilde{\boldsymbol{\theta}}_{i}\right),$ (1)
where $L\colon\mathbb{R}\times\mathbb{R}\to\mathbb{R}_{\geq 0}$ is a scalar loss (e.g., squared or Huber). Following Eq. (1), an estimate of $\boldsymbol{\theta}_{0}$ is obtained as
$\hat{\boldsymbol{\theta}}_{0}=\hat{\theta}_{\mathrm{pre}}(\boldsymbol{z}_{0}),$ (2)
@@ II Problem Statement
When $\boldsymbol{\theta}_{0}\in\Theta_{p}$, suitable choices of $\mathcal{F}$ and $L$ can lead to an accurate estimate $\hat{\theta}_{\mathrm{pre}}(\boldsymbol{z}_{0})$ of $\boldsymbol{\theta}_{0}$. However, because $\Theta_{p}$ is user-specified and may omit plausible regimes, an out-of-distribution (OOD) situation may occur if $\boldsymbol{\theta}_{0}\notin\Theta_{p}$, or if there are not enough samples $\tilde{\boldsymbol{\theta}}_{i}$ in the neighborhood of $\boldsymbol{\theta}_{0}$. Hence, performance can degrade under an OOD, motivating mechanisms that refine $\hat{\theta}_{\mathrm{pre}}(\cdot)$ at implementation phase. In this paper, we focus on a special class of simulation-driven estimators $\hat{\theta}_{\mathrm{pre}}(\cdot)$, known as the Two-Stage (TS) estimator.
II-A The Two-Stage (TS) Estimator
The TS estimator [4, 5] decomposes (1) into two-stages: a fixed compression of $\boldsymbol{z}$ into low-dimensional features, followed by a supervised machine learning model that map these features to the parameters.
A.1 Stage 1 (Compression)
@@
A.2. Stage 2 (Supervised mapping)
@@
A.2.1 Trunk Layers
The trunk of the second stage neural network is mathematically described as follows: Let $\boldsymbol{s}^{(0)}=\boldsymbol{x}$ be the input layer and for $r=1,\dots,R-1$, the output of the $r^{\mathrm{th}}$ hidden layer is given by $\boldsymbol{s}^{(r)}\;=\;\sigma_{r}\!\big(\boldsymbol{W}_{r}\,\boldsymbol{s}^{(r-1)}\big),\,\,\boldsymbol{W}_{r}\in\mathbb{R}^{p_{r}\times p_{r-1}},$ where $\boldsymbol{W}_{r}$ denotes the weights of the hidden layer $r$, $\sigma_{r}(\cdot)$ denotes an activation function (e.g. ReLU($\cdot$)), $p_{0}=n$ and $p_{R-1}=p$. These $R$ layers constitute the trunk of $g_{\mathrm{pre}}$. The trunk output is $\boldsymbol{t}(\boldsymbol{x})\coloneqq\boldsymbol{s}^{(R-1)}\in\mathbb{R}^{p}$ and its parameters are $\boldsymbol{\phi}\coloneqq\{\boldsymbol{W}_{r}\}_{r=1}^{R-1}.$
A.2.2 Head Layers
To finally predict the targets $\tilde{\boldsymbol{\theta}}_{i}\in\Theta_{p}\subseteq\mathbb{R}^{d}$, which can be multidimensional ($d>1$), we append V1V-1 subsequent hidden layers for each parameter dimension, and these V1V-1 layers corresponding to each individual parameter constitute the head of $g_{\mathrm{pre}}(\cdot)$, denoted as $\boldsymbol{h}_{j}^{(v)}$, where $v=1,\ldots,V-1$. We additionally impose that the weights of the first of these V1V-1 layers are least correlated, which is needed to ensure that each parameter is correctly identified. We precisely formalize the structure of the head as the following:
For each parameter index $j\in\{1,\dots,d\}$, starting from $\boldsymbol{h}_{j}^{(0)}=\boldsymbol{t}(\boldsymbol{x})$ and for $v=1,\dots,V-1$ define
$\boldsymbol{h}_{j}^{(v)}\;=\;\sigma_{j,v}\!\big(\boldsymbol{U}_{j,v}\,\boldsymbol{h}_{j}^{(v-1)}\big),\qquad\boldsymbol{U}_{j,v}\in\mathbb{R}^{q_{j,v}\times q_{j,v-1}},$ (5)
where $\boldsymbol{U}_{j,v}$ corresponds to the weights of hidden layer $v$ of head $j$, and $\sigma_{j,v}(\cdot)$ corresponds to the activation function of $v^{\mathrm{th}}$ hidden layer of head $j$, with $q_{j,0}=p$. The scalar output for head $j$ is linear:
$\widehat{{\theta}}_{j}(\boldsymbol{x})\;=\;\boldsymbol{a}_{j}^{\top}\boldsymbol{h}_{j}^{(V-1)},\qquad\boldsymbol{a}_{j}\in\mathbb{R}^{q_{j,V-1}}.$ (6)
The head parameters are then defined as $\boldsymbol{\psi}_{j}~\coloneqq\{(\boldsymbol{U}_{j,v})_{v=1}^{V-1},\boldsymbol{a}_{j}\},$ and $\boldsymbol{\Psi}\coloneqq(\psi_{1},\dots,\psi_{d})$. Stacking the $d$ heads yields the second stage output as
@@ A.2.2 Head Layers
where $\tilde{w}_{j}\geq 0$ are per-parameter weights, $\tilde{{\boldsymbol{\theta}}}_{i}]_{j}$ denotes the $j^{\mathrm{th}}$ coordinate of $\tilde{\boldsymbol{\theta}}_{i}$, $\lambda_{\mathrm{orth}}\geq 0$ is a regularization coefficient, and the last term encourages diversity among the heads by penalizing the correlations between their first-layer weights.
The final TS estimator is the composition $\hat{\theta}_{\mathrm{pre}}(\boldsymbol{z})=g_{\mathrm{pre}}\!\big({h}(\boldsymbol{z});\,\hat{\boldsymbol{\phi}},\hat{\boldsymbol{\Psi}}\big)\in\mathbb{R}^{d}.$ At implementation phase, given a new trajectory $\boldsymbol{z}_{0}$ generated by $\mathcal{M}(\boldsymbol{\theta}_{0})$, the point estimate is
@@
If $\boldsymbol{\theta}_{0}\notin\Theta_{p}$, then $(\boldsymbol{z}_{0},\boldsymbol{\theta}_{0})$ is OOD w.r.t. $\mathcal{D}_{\mathrm{tr}}$, and (9) may yield a poor estimate of $\boldsymbol{\theta}_{0}$. We shall now state a fine-tuning approach to overcome this problem.
III Fine-Tuning the TS Estimator
We propose a simulation-driven transfer learning procedure to improve a pretrained TS estimator when faced with out-of-distribution (OOD) observations. The key idea is to keep the first stage feature extractor $h$ unchanged and to fine-tune only the last layers of the second stage $g_{\mathrm{pre}}(\cdot;\boldsymbol{\phi},\boldsymbol{\Psi})$. Concretely, we freeze the parameters $\boldsymbol{\phi}$ learned during pretraining and update only the remaining parameters $\boldsymbol{\Psi}$ using a small, targeted synthetic dataset produced by the DT in a neighborhood of the current TS estimate. This follows standard transfer-learning practice—earlier layers capture generic structure and are preserved; only the final layers are tuned [20].
Because fine-tuning adds computation and is unnecessary for in-distribution data, we first perform a feature-space statistical test to decide whether an OOD has occurred. If an OOD is detected, we (i) align the simulator’s features to the observed ones via a few Gauss–Newton (GN) steps around the initial TS estimate, and then (ii) use the geometry described by the Fisher information matrix at that estimate to sample an informative cloud of nearby parameters. By using the DT to generate more synthetic data for these samples, we obtain labeled feature–parameter pairs concentrated near the operating point. Finally, we retrain only the final layers of $g_{\mathrm{pre}}$ on this targeted synthetic dataset, keeping $\boldsymbol{\phi}$ fixed (cf. (8)). In summary, the transfer learning approach to fine tune TS consists of the following three steps:

### v3

@@ Abstract
{IEEEkeywords}
Estimation, Identification, Two-Stage Estimator, Machine Learning.
1 Introduction
Recent advances in high-fidelity simulation have popularized digital twins (DTs)–software replicas of physical systems that let engineers explore operating conditions safely and cheaply before deployment [9, 10, 17]. In practice, once a DT is engineered, most modeling choices are fixed; what remains is calibrating a small set of parameters so the DT replicates a specific physical system.
When a DT is available, one can generate synthetic input–output data [15] over a range of parameter settings and train a supervised machine learning model that maps observations to these settings, thereby yielding a simulation-driven estimator. At implementation, this simulation-driven estimator takes real measurements and returns parameter estimates [4, 6, 7]. The Two-Stage (TS) estimator [6] is an attractive simulation-driven estimator with a simple structure that involves two stages: (i) first, the observations are compressed into representative features, and then (ii) these compressed features, along with their corresponding parameter values, are used to train a supervised machine learning model. This structure avoids the need for an explicit likelihood function and it requires low computation time [11, App. B.2] for producing estimates, due to extensive offline training, compared to standard point estimators like the prediction error method (PEM) and the dual Extended Kalman Filter (dual-EKF) [19]. The TS estimator can also be used to learn control design rules, effectively taking the human out of the loop [3].
@@ 1 Introduction
Our novel fine-tuning procedure first compares the compressed features from real data to those produced by the DT at the TS initial prediction, to determine whether an OOD case has occurred by applying a statistical test based on a whitened feature discrepancy measure constructed from estimated feature covariances. If an OOD case is detected, we: (i) compute Gauss–Newton (GN) iterations on the initial prediction by TS, (ii) generate a confidence set around the GN update and re-simulate a targeted local training dataset with the DT, and (iii) fine-tune the final portion of the second stage of the TS estimator, which is inspired from transfer learning [14] approaches to system identification [5, 13]. Our work differs from modern simulation-based inference methods [24, 2, 20], which aim to infer a full Bayesian posterior. In addition, within these methods, some neural posterior estimation methods use differentiable simulator gradients [24], and others use truncated proposals for sequential refinement [2]. In contrast, our approach is a non-Bayesian scheme for a pre-trained point estimator, requiring no differentiable simulator. We perform a lightweight, final-layer update, which is conceptually closer to other transfer learning methods in system identification [5, 13] and broader OOD adaptation strategies [12, Ch. 19-20]. In particular, our contributions are the following:
@@
The remainder of the paper is organized as follows. In Section 2, we introduce the problem statement. Section 3 describes the TS estimator. In Section 4, our method for fine-tuning the pretrained TS estimator is detailed, and in Section 5, we demonstrate the effectiveness of our approach through several numerical examples. Section 6 outlines the future work. Finally, Section 7 concludes the paper.
2 Problem Statement
We consider a data-generating mechanism $\mathcal{M}(\boldsymbol{\theta})$, a high-fidelity simulator or DT, parameterized by $\boldsymbol{\theta}\in\Theta\subseteq\mathbb{R}^{d}$. For notational simplicity, we assume a single-input single-output (SISO) setting: given an $N$-length scalar input sequence $\boldsymbol{u}=(u_{1},\dots,u_{N})^{\top}\in\mathbb{R}^{N}$, the DT produces a (possibly stochastic) scalar output sequence $\boldsymbol{y}=(y_{1},\ldots,y_{N})^{\top}\in\mathbb{R}^{N}$. Concretely, $\boldsymbol{y}=\boldsymbol{y}(\boldsymbol{\theta};\omega)$, where $\omega$ denotes the simulator’s noise seed (or any internal randomness). Assume that a length-$N$ trajectory $\boldsymbol{z}_{0}=(u_{1},y_{0,1},\ldots,u_{N},y_{0,N})\in\mathbb{R}^{2N}$ has been measured from the DT at some unknown parameter $\boldsymbol{\theta}_{0}\in\Theta$. The goal is to estimate $\boldsymbol{\theta}_{0}$ from $\boldsymbol{z}_{0}$ by constructing a simulation-driven parameter estimator.
Given a calibration set $\Theta_{p}\subseteq\Theta$, constructed from prior domain knowledge and possible operating regimes, we draw independent and identically distributed (i.i.d.) parameter settings $\{\tilde{\boldsymbol{\theta}}_{i}\}_{i=1}^{m}$ from $\Theta_{p}$. For each draw $\tilde{\boldsymbol{\theta}}_{i}$, we run the simulator with (possibly) its own noise seed $\omega_{i}$ and the chosen input sequence [^1] Due to space limitations, in this paper the input sequence is considered fixed, even though the treatment can be extended to random inputs. $\boldsymbol{u}$, obtaining $\boldsymbol{z}_{i}=(u_{1},y_{i,1},\ldots,u_{N},y_{i,N})$. This yields a synthetic, labeled dataset $\mathcal{D}=\{(\boldsymbol{z}_{i},\tilde{\boldsymbol{\theta}}_{i})\}_{i=1}^{m}$. An estimator $\hat{f}_{\mathrm{pre}}\colon\mathbb{R}^{2N}\!\to\mathbb{R}^{d}$ is then fitted by solving the following minimization problem over a prescribed function class $\mathcal{F}$ (e.g. a neural network):
$\hat{f}_{\mathrm{pre}}\in\arg\min_{F\in\mathcal{F}}\frac{1}{m}\sum_{i=1}^{m}L\!\left(F(\boldsymbol{z}_{i}),\,\tilde{\boldsymbol{\theta}}_{i}\right),$ (1)
where $L\colon\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}_{\geq 0}$ is a suitable loss function. Following Eq. (1), an estimate of $\boldsymbol{\theta}_{0}$ is obtained as
$\hat{\boldsymbol{\theta}}_{0}=\hat{f}_{\mathrm{pre}}(\boldsymbol{z}_{0}),$ (2)
@@ 2 Problem Statement
When $\boldsymbol{\theta}_{0}\in\Theta_{p}$, suitable choices of $\mathcal{F}$ and $L$ can lead to an accurate estimate $\hat{f}_{\mathrm{pre}}(\boldsymbol{z}_{0})$ of $\boldsymbol{\theta}_{0}$. However, because $\Theta_{p}$ is user-specified and may omit plausible regimes, an out-of-distribution (OOD) situation may occur if $\boldsymbol{\theta}_{0}\notin\Theta_{p}$, or if there are not enough samples $\tilde{\boldsymbol{\theta}}_{i}$ in the neighborhood of $\boldsymbol{\theta}_{0}$. Hence, performance can degrade under an OOD, motivating mechanisms that refine $\hat{f}_{\mathrm{pre}}$ at implementation phase. In this paper, we focus on a special class of simulation-driven estimators $\hat{f}_{\mathrm{pre}}$, known as the TS estimator.
3 The TS Estimator
The TS estimator [6] decomposes (1) into two-stages: a fixed compression of $\boldsymbol{z}$ into low-dimensional features, followed by a supervised machine learning model that map these features to the parameters.
3.1 Stage 1 (Compression)
@@
3.2 Stage 2 (Supervised mapping)
@@
3.2.1 Trunk Layers
The trunk of the second stage neural network is mathematically described as follows: Let $\boldsymbol{c}^{(0)}=\boldsymbol{x}$ be the input layer and for $r=1,\dots,R-1$, the output of the $r^{\mathrm{th}}$ hidden layer is given by $\boldsymbol{c}^{(r)}\;=\;\sigma_{r}\!\big(\boldsymbol{W}_{r}\,\boldsymbol{c}^{(r-1)}\big),\,\,\boldsymbol{W}_{r}\in\mathbb{R}^{p_{r}\times p_{r-1}},$ where $\boldsymbol{W}_{r}$ denotes the weights of the hidden layer $r$, $\sigma_{r}$ denotes an activation function (e.g. ReLU($\cdot$)), $p_{0}=n$ and $p_{R-1}=p$. These $R$ layers constitute the trunk of $g_{\mathrm{pre}}$. The trunk output is $\boldsymbol{t}(\boldsymbol{x})\coloneqq\boldsymbol{c}^{(R-1)}\in\mathbb{R}^{p}$ and its parameters are $\boldsymbol{\phi}\coloneqq\{\boldsymbol{W}_{r}\}_{r=1}^{R-1}.$
3.2.2 Head Layers
To finally predict the targets $\tilde{\boldsymbol{\theta}}_{i}\in\Theta_{p}\subseteq\mathbb{R}^{d}$, which can be multidimensional ($d>1$), we append V1V-1 subsequent hidden layers for each parameter dimension, and these V1V-1 layers corresponding to each individual parameter constitute the head of $g_{\mathrm{pre}}$, denoted as $\boldsymbol{h}_{j}^{(v)}$, where $v=1,\ldots,V-1$. We additionally impose that the weights of the first of these V1V-1 layers are least correlated, which is needed to ensure that each parameter is correctly identified. We precisely formalize the structure of the head as the following:
For each parameter index $j\in\{1,\dots,d\}$, starting from $\boldsymbol{h}_{j}^{(0)}=\boldsymbol{t}(\boldsymbol{x})$ and for $v=1,\dots,V-1$ define $\boldsymbol{h}_{j}^{(v)}\;=\;\sigma_{j,v}\!\big(\boldsymbol{U}_{j,v}\,\boldsymbol{h}_{j}^{(v-1)}\big),\,\boldsymbol{U}_{j,v}\in\mathbb{R}^{q_{j,v}\times q_{j,v-1}},$ where $\boldsymbol{U}_{j,v}$ corresponds to the weights of hidden layer $v$ of head $j$, and $\sigma_{j,v}$ corresponds to the activation function of $v^{\mathrm{th}}$ hidden layer of head $j$, with $q_{j,0}=p$. The scalar output for head $j$ is linear: $\widehat{{\theta}}_{j}(\boldsymbol{x})\;=\;\boldsymbol{a}_{j}^{\top}\boldsymbol{h}_{j}^{(V-1)},\qquad\boldsymbol{a}_{j}\in\mathbb{R}^{q_{j,V-1}}.$ The head parameters are then defined as $\boldsymbol{\psi}_{j}~\coloneqq\{(\boldsymbol{U}_{j,v})_{v=1}^{V-1},\boldsymbol{a}_{j}\},$ and $\boldsymbol{\Psi}\coloneqq(\psi_{1},\dots,\psi_{d})$. Stacking the $d$ heads yields the second stage output as
@@ 3.2.2 Head Layers
where $\tilde{w}_{j}\geq 0$ are weights chosen as $\tilde{w}_{j}\;=\;1/{\widehat{\mathrm{Var}}([\tilde{\boldsymbol{\theta}}]_{j})}$, where ${\widehat{\mathrm{Var}}([\tilde{\boldsymbol{\theta}}]_{j})}\;:=\;m^{-1}\sum_{i=1}^{m}([\tilde{\boldsymbol{\theta}}_{i}]_{j}-\overline{\theta}_{j})^{2},\,\overline{\theta}_{j}:=m^{-1}\sum_{i=1}^{m}[\tilde{\boldsymbol{\theta}}_{i}]_{j}.$ Here, $l$ is the scalar Huber loss [1], $\tilde{{\boldsymbol{\theta}}}_{i}]_{j}$ denotes the $j^{\mathrm{th}}$ coordinate of $\tilde{\boldsymbol{\theta}}_{i}$, $\lambda_{\mathrm{orth}}\geq 0$ is a regularization coefficient, and the last term encourages diversity among the heads by penalizing the correlations between their first-layer weights. Notice that the overall loss function $L$ in Eq. (1) is related to $l$ as $L(\hat{\boldsymbol{\theta}},\boldsymbol{\theta}):=\sum_{j=1}^{d}\tilde{w}_{j}\,\ell([\hat{\boldsymbol{\theta}}]_{j},\,[\boldsymbol{\theta}]_{j})+\lambda_{\mathrm{orth}}\!\sum_{1\leq j<k\leq d}\!\big\|\boldsymbol{U}_{j,1}\boldsymbol{U}_{k,1}^{\top}\big\|_{F}^{2}.$
The final TS estimator is the composition $\hat{f}_{\mathrm{pre}}(\boldsymbol{z})=g_{\mathrm{pre}}\!\big({h}(\boldsymbol{z});\,\hat{\boldsymbol{\phi}},\hat{\boldsymbol{\Psi}}\big)\in\mathbb{R}^{d}.$ At implementation phase, given a new trajectory $\boldsymbol{z}_{0}$ generated by $\mathcal{M}(\boldsymbol{\theta}_{0})$, the estimate is
@@
If $\boldsymbol{\theta}_{0}\notin\Theta_{p}$, then $(\boldsymbol{z}_{0},\boldsymbol{\theta}_{0})$ is OOD w.r.t. $\mathcal{D}_{\mathrm{tr}}$, and Eq. (7) may yield a poor estimate of $\boldsymbol{\theta}_{0}$. We shall now state a fine-tuning approach to overcome this problem.
4 Fine-Tuning the TS Estimator
We propose a simulation-driven transfer learning procedure to improve a pretrained TS estimator under OOD observations. We keep the first stage feature extractor $h$ unchanged and fine-tune only the last layers of the second stage $g_{\mathrm{pre}}(\cdot;\boldsymbol{\phi},\boldsymbol{\Psi})$: we freeze the pretrained $\boldsymbol{\phi}$ and update only the head layers $\boldsymbol{\Psi}$ using a small, targeted synthetic dataset from the DT in a neighborhood of the current TS estimate, following standard transfer learning practice i.e., earlier layers capture generic structure and are preserved, so only the final layers are tuned [23]. In summary, the procedure consists of the following steps: